E, (V/m) 29. What is the potential difference V-V, between Xi = 0 m and x,-1.0 m? 200 B. 50 V C.-50 V D. 100V E.-100 V F. 150 V G.-150V 100 x (m) -100
Answer #1

to find the potential diffrence we find area of triangle,

area of triangle = (1/2)*height*base

area of triangle = 1/2 * 100*1 = 50

so Vf -Vi = 50V